OpenStudy (anonymous):

What is the slope of the line that passes through the given points (3,2)and(5,12) answers are 5. -5. -1/5. 1/5. Thanks

OpenStudy (texaschic101):

use the slope formula... slope = (y2 - y1) / (x2 - x1) (3,2) x1 = 3 and y1 = 2 (5,12) x2 = 5 and y2 = 12 now sub slope = (12 - 2) / (5 - 3) slope = 10/2 slope = 5